OpenStudy (anonymous):

) Jack spends 1 1/2 as long on his homework as Jill. Lask week, Jill spent 8 1/2 hours doing homework. How long did Jack spend doing homework?

OpenStudy (alekos):

so that would be 1&1/2 x 8&1/2 hrs

OpenStudy (jack1):

so whats up with laTex on this site at the mo'?

OpenStudy (anonymous):

\[8\frac{ 1 }{ 2 }\times.50=4\frac{ 1 }{ 4 }\]\[4\frac{ 1 }{ 4 }+8\frac{ 1 }{ 2 }=12\frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

So the answer is, he spent the whole day doing homework. He spent:\[12\frac{ 3 }{ 4 }\]hours on his homework.

OpenStudy (anonymous):

@aliyaaleena

OpenStudy (anonymous):

You could use both alekos's way, or mine. Both come out to the same answer.